OpenStudy (anonymous):

Solve the equations: (4x-6)(2x-5)=0

OpenStudy (anonymous):

x=6/4=3/2 or x=5/2

OpenStudy (anonymous):

i think i solved this earlier..

OpenStudy (anonymous):

okaeeee . and how would i put it in brackets? and@ arnab09 you did but i wrote the problem wrong :)

OpenStudy (anonymous):

{3/2,5/2}

OpenStudy (anonymous):

order doesnot matter:)

OpenStudy (anonymous):

Then how would i do this one? 3x^2=6 ?

OpenStudy (anonymous):

With brackets i think .

OpenStudy (anonymous):

sqart(2, -2)

OpenStudy (anonymous):

\[x=\pm \sqrt{2}\]

OpenStudy (anonymous):

3x^2-6=o 3(x^2-2)=0 x^2-2=0 then x=\[\pm{\sqrt{2}}\]

OpenStudy (anonymous):

@ koena, bracket is not applicable all time.. :)

OpenStudy (anonymous):

Ok :)

OpenStudy (anonymous):

Thanks everyone that really helped!